Minutes PSB Upgrade WG Meeting 198, 1st of August 2017

Participants: Simon Albright, Michele Battistin, Jean-Baptiste Bonnamy, Jean-Pierre Corso, James Dilwyn Devine, Jose Antonio Ferreira Somoza, Alessandro Floriduz, Eve Fortescue-Beck, Robert Foreschl, Georgi Minchev Georgiev, Greta Guidoboni, Klaus Hanke, David Hay, Bettina Mikulec, Antony Newborough, Jérôme Rodary, Andrea Santamaria Garcia, Wim Weterings.

Indico link:

1. Approval of Minutes

The minutes of the LIU-PSB WG #197 were approved.

2. Communications

  • LIU Project Team meeting on the 20th of July 2017.
  • The budget for the Linac 2 dismantling is currently under discussion. There are 25 groups involved in this operation, and the budget will be split between the main parties.

3. Follow-up of Open Actions

Actions to be followed-up after holidays (change of deadline to the end of September):

  • J-M. Lacroix (2017-06-15): Complete integration of PSB-PS transfer line.
  • B. Riffaud (2017-06-15): Provide a proposal for the support of the equipment in the BTP line.

New actions (the deadline will be postponed until the 29th of August):

  • T. Dobers (2017-08-01): Make a list of starting points of the transfer lines to be sent for approval in an EDMS document.

Related to the technical topic (the data needs to be cross-checked with the users, so the action will be delayed until the end of the summer):

  • G. Petrika (2017-08-01): Prepare a document summarizing the ventilation needs for LIU-PSB

Update on the PSB Ventilation and Cooling Systems by Gabor Petrika (slides)

  • The user requirements for cooling are summarized in the EDMS document 1758628.
  • The 3 Finemet cavities are critical user locations for cooling, so new dedicated branches will be built for them. To this end two new pipes will be installed, in addition to the existing ones, that will go from the surface to the critical users.
  • The first version of the P&ID for cooling is foreseen for December 2017.
  • The mean heat load per rack was measured on-site with EN-EL and EN-CV.
  • Two ventilation systems will exist: one for the tunnel and one for the surface. A new operational mode, the foil change mode, has been added to the tunnel ventilation system.
  • It is foreseen to reduce the work load in LS2 by installing some parts beforehand, namely the smoke extraction system and ventilation, the water treatment system, the piping gallery and the piping instrumentation.
  • During LS2 the ventilation system cannot be switched-off, so basic ventilation will be available.
  • By November 2019 the RF systems will need cooling, so the system must be available for the users by that time.


  • B. Mikulec suggested to cross-check the room temperature value (25 ± 5 °C) with Alan Findlay too see if it meets the requirements of the low-level RF since it might need to be lower in some areas.
  • B. Mikulec asked for a more precise estimation of the time of the flush mode, which is currently estimated to be of half an hour.
  • B. Mikulec asked if there will be spare parts to repair the ventilation system. It needs to be understood which equipment will trip at a certain temperature.
  • B. Mikulec asked to know more details on how the different ventilation modes can be controlled from the CCC with respect to the access system.
  • D. Hay asked if the terminal heaters could be removed during the YETS instead of LS2, since it’s currently unused equipment. P. Gabor answered that the tunnel ceiling is 5 meters tall and that a temporary structure must be built for that task. Additionally, one has to tackle the issues related to the dismantling of activated units and the potential presence of asbestos. He agreed to ask the contractor for a time estimate for this task, and see if it’s possible to do it during the YETS (a VIC will be done during the September technical stop). J. Devine commented that the spare place left after the removal of the units would be the perfect place to install new lighting.
  • It was agreed that the installation of the new piping will be cross-checked with J-P. Corso from integration, once all the details are known. J-P. Corso commented that there might be a space constraint to accommodate the new pipes in LS2.  
  • P. Gabor commented that he is currently working on the budget.

5. AOB

Concerning the tour de table:

  • J. Devine reported that the document summarizing the lighting needs for building 361 and the surrounding buildings is ready for EL approval.
  • R. Froeschl commented that the first measurements regarding the removal works of activated material will be done during the next YETS.
  • A. Newborough communicated that the price of the recombination dipoles and correctors is 3 times higher than estimated. They are now looking into different solutions and will report back on the topic.


Assigned to Due date Description State Result  
K.Hanke 2017-09-30 Discuss with SMB the consolidation of building 141 edit
G.Petrika 2017-12-05 Iterate with integration on the installation of the new piping system for cooling. edit
G.Petrika 2017-10-10 release a new version of the EDMS document 1758628, summarizing the cooling needs for LIU-PSB. edit
G.Petrika 2017-09-30 Check the viability of the removal of the obsolete heater units during the YETS during TS2. edit
A.Newborough 2017-09-30 Provide the final price of the recombination dipoles and correctors. edit

LIU-PSB meeting on the 15th of August 2017: canceled.

Next LIU-PSB meeting: 29th of August.

-- AndreaSantamariaGarcia - 2017-08-07

Edit | Attach | Watch | Print version | History: r7 < r6 < r5 < r4 < r3 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r7 - 2018-01-10 - unknown
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSBUpgrade All webs login

This site is powered by the TWiki collaboration platform Powered by PerlWeb page setup by Thomas Hermanns